Y= -2 x
-------
3
2x + 3y = 5

Answer by Mathtut(3670)

y=-2x+3--->2x+y=3..eq 1
2x+3y=5...eq2
:
subtract eq 2 from eq 1...x terms are eliminated. we are left with y-3y=3-5
:
-2y=-2
:
y=1
:
now take y's found value and plug it back into either equation. I choose eq 2
:
2x+3(1)=5
:
2x=2
:
x=1
:
(1,1)is the solution